A particle moves on a line with a constant acceleration of 4ft/second. Find its equation of motion if s=2ft and v=-ft/sec, when t=2sec


1
Expert's answer
2022-01-27T13:16:48-0500

Newton motion second law

"s=ut+\frac{1}{2}at^2"

"2=4\times2+\frac{1}{2}\times a\times2^2"

Solve it

"2-8=\frac{1}{2}a\times2^2"

"-6\times2=2^2a"

"a=-3ft\/sec^2"

Where

"v=4ft\/sec"

"s=2ft\\u=4ft/sec\\a=-3ft/sec^2"